ure of the object itself should not change. However, if a significant amount of heat flows from the object to speche heat capacey ofc-81SJ/(kg·C), and a temperature of 12.0 С It is immersed ture will change. A thermometer has a mass of 36.0g, a and the final temperature of the water and thermometer is 41.29 C. What was the temperature of the water before t in 119 g of water, the insertion of the thermometer? (See Table 12.2 for appropriate constants.)
Answer #1

Given : Mt = 36 g or 0.036 kg , Ct = 815 J/kg°c , Tt = 12°c , Mw= 119 g or 0.119 kg , Tf = 41.2°c

Goal : Initial temperature of water (Tw)= ?

Constant : heat capacity of water (Cw)= 4186 J/kg°c

Solution:

According to the conservation of energy,

Heat loss by water = Heat gain by thermometer

MwCw(Tw-Tf) = MtCt(Tf-Tt)

(0.119)(4186)(Tw-41.2) = (0.036)(815)(41.2 - 12)

498.134(Tw-41.2) = 856.728

Tw - 41.2 = 1.72

Tw = 42.92°c

Answer : Tw = 42.92°c